Suppose that the relationship between te tempeature t of a cup of water (in degrees fahrenheit) and the number of seconds s that is in themicrowave is given by linear equation. If initially, the temperature is 60 degrees fahrenheit, and after 120 seconds it has risen to 200 degrees fahrnheit:
1) What would the temperature be after 30 seconds?
2) How long would youhave to leave it in for the temperature to reach 300 degrees fahrenheit?

Answer by stanbon(75887)

Suppose that the relationship between te tempeature t of a cup of water (in degrees fahrenheit) and the number of seconds s that is in themicrowave is given by linear equation. If initially, the temperature is 60 degrees fahrenheit, and after 120 seconds it has risen to 200 degrees fahrnheit:
-----------
You are given two "points" relating time and temperature: (0,60), (120,200)
From these you can form an equation:
slope = [200-60]/[120-0]= 140/120=7/6
y-60=(7/6)(x-0)
y=(7/6)x+60
Y is temperature and x is time.
1) What would the temperature be after 30 seconds?
temp = (7/6)(30)+60 = 95 degrees
-------------
2) How long would you have to leave it in for the temperature to reach 300 degrees fahrenheit
300=(7/6)x+60
240=(7/6)x
x=(6/7)240
time= 205.71 seconds
--------
Cheers,
Stan H.